OpenStudy (anonymous):

rewrite the addition sum given below using digits. Each letter is used for a different digit. Answer the questions given below i) Which digit did you find first? ii) How did you find R and U? ........F.O.U.R ...........O.N.E ....T.H.R.E.E ....T.H.R.E.E _________ E.L.E.V.E.N

OpenStudy (shubhamsrg):

four = 1000f + 100o + 10u +r one= 100o + 10n +e three= 10000t + 1000h + 100r + 10e + e eleven = 100000e + 10000l +1000e +100v + 10e + n simplify and see if it helps.